OpenStudy (anonymous):

How would you factor this.? Steps please. 2x^2=-24x-72

OpenStudy (anonymous):

First you want to get all terms on the left side and set it equal to zero. Do you know how to do that?

OpenStudy (anonymous):

\[2x^2 + 24x + 72 = -24x + 24x -72 + 72\] Add 24x and 72 to both sides of the equation.

OpenStudy (anonymous):

\[2x^2 + 24x + 72 = 0\]

OpenStudy (jdoe0001):

\(\large {2x^2=-24x-72\implies 2x^2+24x+72=0 \\ \quad \\ \implies \begin{array}{cccllll} 2(x^2&+12x&+36)=0\\ &\uparrow &\uparrow\quad \\ &6+6&6\cdot 6 \end{array}}\)

OpenStudy (anonymous):

\[2(x^2 + 12x + 36) = 0\]

OpenStudy (anonymous):

\[\frac{2}{2}(x^2 + 12x + 36) = \frac{0}{2}\]

OpenStudy (anonymous):

Divide both sides by 2 to eliminate the 2 on the left side of the equation. From there you factor using the description @jdoe0001 gave three posts up.

OpenStudy (anonymous):

You are looking for "what terms multiply to get 36 and add to get 12?"

OpenStudy (anonymous):

Since all the terms of your polynomial are positive, the terms of your factored form will all be positive.

OpenStudy (anonymous):

\[(x + \_\_)(x+\_\_) = 0\]

OpenStudy (anonymous):

Fill in the blanks.